# Hybrid Monte Carlo for Failure Probability Estimation with Gaussian Process Surrogates

### v1

@@ Abstract
We tackle the problem of quantifying failure probabilities for expensive computer experiments with stochastic inputs. The computational cost of evaluating the computer simulation prohibits direct Monte Carlo (MC) and necessitates a statistical surrogate model. Surrogate-informed importance sampling – which leverages the surrogate to identify suspected failures, fits a bias distribution to these locations, then calculates failure probabilities using a weighted average – is popular, but it is data hungry and can provide erroneous results when budgets are limited. Instead, we propose a hybrid MC scheme which first uses the uncertainty quantification (UQ) of a Gaussian process (GP) surrogate to identify areas of high classification uncertainty, then combines surrogate predictions in certain regions with true simulator evaluation in uncertain regions. We also develop a stopping criterion which informs the allocation of a fixed budget of simulator evaluations between surrogate training and failure probability estimation. Our method is agnostic to surrogate choice (as long as UQ is provided); we showcase functionality with both GPs and deep GPs. It is also agnostic to design choices; we deploy contour locating sequential designs throughout. With these tools, we are able to effectively estimate small failure probabilities with only hundreds of simulator evaluations. We validate our method on a variety of synthetic benchmarks before deploying it on an expensive computer experiment of fluid flow around an airfoil.
Keywords: computer experiment, deep Gaussian process, entropy, emulator, contour location, reliability, importance sampling
@@ 1 Introduction
Quantifying the probability of failure in complex dynamical systems is of great interest but is challenging due to the computational expense of the so-called “computer experiment.” Consider a black-box computer simulation, $f:\mathcal{X}\rightarrow\mathbb{R}$ with inputs $\mathcal{X}\subset\mathbb{R}^{d}$ governed by known measure $\mathbb{P}_{\mathbf{x}}$ (with density $p(\mathbf{x})$). The response is thresholded such that $f(\mathbf{x})>t$ indicates failure. The research objective is to estimate the failure probability,
@@
We are particularly motivated by aeronautic applications where system failures may incur immense financial or environmental costs, or even loss of life. In aircraft design and development, computer simulation experiments are a natural, often mandatory, precursor to direct experimentation and deployment. For example, $f(\cdot)$ may represent the vibrations experienced by an aircraft which become unsafe above certain levels, or the fuel efficiency of an aircraft, which if too low, can be detrimental to the environment, obliterate fuel budgets, or overly restrict travel distances. Inputs $\mathcal{X}$ may encapsulate design parameters and operating conditions (such as wing shapes, aircraft design, flight speeds, and angles of attack). Yet we acknowledge a slew of other relevant applications where quantifying failure probabilites for complex simulations is critical, including structural reliability analysis using finite element simulations (e.g., Mia et al.,, 2017), Boltzmann modeling of chemical processes in fuel filtration (e.g., Belot et al.,, 2021), and thermodynamic flows for nuclear reactors (e.g., Batet et al.,, 2014).
@@
A rudimentary approach to solve Eq. (1) involves brute force Monte Carlo (MC); simply sample an exhaustively large set of inputs, evaluate the computer simulation at each of these configurations, and report the proportion that resulted in failure. But this approach is prohibitively expensive if the computer simulation is costly (incurring long wall clock times), the probability of failure is rare (requiring enourmous sample sizes), or both! In this work, we propose a novel method for quantifying failure probabilities of complex computer simulations when the computation budget restricts evaluations of the simulator to several hundred, rendering traditional MC impractical.
@@
In the face of expensive data, statistical models (called “surrogates” or “emulators”) are essential. Surrogates are trained on a limited set of simulator evaluations in order to provide predictions with appropriate uncertainty quantification (UQ) at unobserved inputs. Gaussian processes (GPs) are the canonical choice (Santner et al.,, 2003; Gramacy,, 2020), but deep Gaussian processes (DGPs; Damianou and Lawrence,, 2013) are becoming increasingly popular, particularly for non-stationary computer experiments (Sauer,, 2023). Notably, our contribution is agnostic to surrogate choice as long as effective UQ is provided, and is thus applicable to traditional GPs and Bayesian DGPs (Sauer et al.,, 2023b) and, potentially, Bayesian neural networks (MacKay,, 1995).
@@
Early methods of failure probability estimation, namely the first and second order reliability methods (FORM/SORM; Haldar and Mahadevan,, 1995), leveraged crude linear approximations of the complex model and placed all failure probability estimation in the hands of the “most probable point” on the predicted contour. Advancements on FORM and SORM have been made with more flexible GP models (e.g., Su et al.,, 2017). A variety of works have instead leveraged GP surrogates directly with MC sampling, using the GP predictions (which are of neglible cost) in place of true simulator evaluation (e.g., Cheng and Lu,, 2020; Li et al.,, 2021; Lu et al.,, 2023). We will use this “surrogate MC” method as a benchmark in our later exercises. While it is cheap, it places too much stock in the accuracy of the surrogate and funnels surrogate inaccuracies directly into the failure probability estimate.
Importance sampling (IS; Srinivasan,, 2002; Tokdar and Kass,, 2010) offers an unbiased alternative. It hinges on the construction of a bias distribution that intentionally focuses density in the failure region. Inputs are sampled from the bias distribution (hopefully yielding many failures) and are then utilized in a weighted average calculation that corrects for the disparity between the bias density and the original input density. Identifying an effective bias distribution in the face of limited training data is a hefty task (Tabandeh et al.,, 2022). Adaptive methods abound (e.g., Oh and Berger,, 1992; Kurtz and Song,, 2013; Dalbey and Swiler,, 2014; Bugallo et al.,, 2017; Cheng et al.,, 2023). Peherstorfer et al., (2016) propose what they termed “multifidelity importance sampling,” although we prefer the term “surrogate-informed importance sampling” (SIIS) to avoid confusion with multifidelity surrogate modeling (e.g., Park et al.,, 2017). In SIIS, a surrogate is used to identify predicted failures, and a bias distribution is fit to these locations (more on this in Section 2.2). SIIS has been deployed with GP surrogates in a variety of subsequent works (e.g., Cole et al.,, 2023; Renganathan et al.,, 2023; Booth et al.,, 2024a; Renganathan,, 2024), and we view it as our primary competitor. Yet SIIS has several crucial drawbacks. We will delay a thorough discussion to Section 2.2, but the punchline is simply that IS has too much variability when bias distributions are suboptimal and data sizes are limited. The difficulty in training an effective bias distribution can render SIIS estimated failure probabilities ineffective, forming the motivation behind our current contribution.
We argue that utility and precision in practice are more important than theoretical unbiasedness. To this end, we embrace a hybrid Monte Carlo calculation that uses both surrogate predictions and true simulator evaluations in strategically partitioned regions of the input space. Li and Xiu, (2010) first proposed such a hybrid calculation, but their work was restricted to polynomial chaos (PC) models with strict accuracy assumptions in order to guarantee convergence. GP-based surrogates afford several key advancements over PCs. For one, GPs are flexible and data-driven; their accuracy continuously improves as training data sizes increase. Perhaps more importantly, GPs offer UQ through posterior variances. We leverage this UQ to prioritize expensive model evaluation at locations where the surrogate is least certain in its pass/fail prediction.
@@
Regardless of how a surrogate is to be used (surrogate MC, SIIS, our hybrid MC, etc.), surrogate accuracy in identifying failure regions is of primary importance. Not all training designs are created equally. Space-filling designs (Joseph,, 2016) may miss failure regions altogether, particularly when training budgets are limited. Instead, it is common to train a surrogate on a small initial design then pair it with a relevant acquisition function to select subsequent inputs for evaluation through the simulator. This iterative process of acquisition and surrogate updating can strategically hone in on the contour, all while minimizing evaluations of the expensive simulator. So-called “contour location” (CL) naturally outperforms space-filling counterparts. Acquisition functions seek to balance exploitation and exploration by leveraging both predicted distance from the contour and posterior uncertainty. Popular variations fall under expected improvement frameworks (Bichon et al.,, 2008; Ranjan et al.,, 2008), stepwise uncertainty reduction frameworks (Bect et al.,, 2012; Chevalier et al.,, 2014; Azzimonti et al.,, 2021), or both (Duhamel et al.,, 2023). Others utilize pass/fail predictive entropy (Marques et al.,, 2018; Cole et al.,, 2023). Booth et al., (2024b) recently extended CL sequential designs to DGP surrogates. We will utilize state-of-the-art sequentially designed GP and DGP surrogates in our benchmark exercises, but our contribution is agnostic to these choices. We see our method’s “plug-and-play” nature as a distinct advantage.
Both surrogate training and failure probability estimation require a budget of expensive simulator evaluations. These two tasks are typically conducted in independent stages, with separate pre-determined budgets (as we will use in the first round of our synthetic benchmarks). Yet, in practice, it is common for a total budget to be set according to computational allowances, with no mention of how many evaluations to spend on the training phase versus the estimation phase. Some works have investigated stopping criteria for surrogate training in reliability settings (e.g., Echard et al.,, 2011; Hristov et al.,, 2019), but this is still an under-examined area of the literature. In our hybrid MC set-up, we have a unique advantage: the more we train our surrogate, the more certain it will be, requiring fewer high-cost evaluations for effective failure probability estimation. There is a clear give-and-take between the budgets of the training and estimation stages; increasing the budget of one decreases the budget required for the other, and vice versa. Leveraging this trade-off, we propose a method for allocating a fixed budget between surrogate training and hybrid Monte Carlo estimation.
This manuscript is laid out as follows. First, in Section 2, we review GP/DGP surrogates and competing methods of failure probability estimation. In Section 3 we present our proposed hybrid Monte Carlo procedure, including implementation details and a variety of synthetic benchmarking exercises. In Section 4 we propose our budget allocation scheme, continuing the established benchmarks. Finally, we deploy our method on a computer experiment of an RAE-8288 airfoil in Section 5 and conclude in Section 6.
2 Failure probability estimation
@@
This estimator is ineffective unless a sufficient number of failures are observed, necessitating very large $M$ particularly for small $\alpha$. But our computer model $f(\mathbf{x})$ is expensive; we simply cannot evaluate it thousands/millions/billions of times. We will describe our sample efficient method for estimating failure probabilities in Section 3, after a quick review of preliminaries.
@@ 2.1 Gaussian process surrogates
A surrogate model $\hat{f}$, trained on a limited budget of runs, may provide predictions at unobserved inputs in place of the true $f$. To enable our proposed methodology the surrogate must also provide posterior predictive distributions. Gaussian processes (GPs) are the predominant choice; they assume a multivariate normal prior distribution over the response, $y\sim\mathcal{N}(\mu,\Sigma(X))$. The prior mean $\mu$ is often simplified to the zero vector (which we will use moving forward), and the prior covariance $\Sigma(X)$ is typically a function of inverse euclidean distances, e.g. $\Sigma(X)^{ij}=\Sigma(\mathbf{x}_{i},\mathbf{x}_{j})=k(||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2})$ with kernel $k(\cdot)$. See Santner et al., (2003); Rasmussen et al., (2006); Gramacy, (2020) for thorough reviews. Conditioned on training data $\left\{X_{n},\mathbf{y}_{n}\right\}$, posterior predictions at singeton input $\mathbf{x}$ follow
$\hat{f}_{n}(\mathbf{x})\sim\mathcal{N}\left(\mu_{n}(\mathbf{x}),\sigma_{n}^{2}(\mathbf{x})\right)\quad\textrm{for}\quad\begin{array}[]{rl}\mu_{n}(\mathbf{x})&=\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\mathbf{y}_{n}\\ \sigma_{n}^{2}(\mathbf{x})&=\Sigma(\mathbf{x})-\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\Sigma(X_{n},\mathbf{x}),\end{array}$ (3)
@@
A natural alternative to true MC (Eq. 2) is “surrogate MC” which leverages surrogate predictions in place of the true simulator,
$\hat{\alpha}_{\textrm{SURR}}=\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}_{\left\{\mu_{n}(\mathbf{x}_{i})>t\right\}}\quad\textrm{for}\quad\mathbf{x}_{i}\sim p(\mathbf{x}).$ (4)
If the surrogate is not accurate near the contour, it will misclassify inputs which will severely hinder the effectiveness of this estimator.
Figure 1: Left/Center: Herbie function with failure contour (red) and input distribution (black). Right: Entropy surface (yellow/high, purple/low) for a GP trained on a 100 point LHS (white circles).
Figure 2: Slice of Herbie function along $x_{2}=1.1$. GP was trained on 100 point LHS in 2 dimensions.
For example, consider the two-dimensional “Herbie” function (Lee et al.,, 2011) shown in the left and center panels of Figure 1 (we will revisit the right panel momentarily). We define a failure threshold of $t=1.065$ which demarcates four disjoint failure regions (red contours). We also define $p(\mathbf{x})$ using independent normal distributions centered at the origin (solid black contours). [^1] Full details of all simulated functions, thresholds, and distributions are provided in the Supplementary Material. We trained a GP surrogate using a random Latin hypercube sample (LHS; McKay et al.,, 2000) of size $n=100$ in the two dimensions. Figure 2 shows the GP posterior prediction along the slice $x_{2}=1.1$. The posterior mean (blue dashed) and 95% CI (blue shading) capture the true nonlinear surface (solid black) fairly well. Even so, there are inputs (highlighted by the yellow shading) which the surrogate would misclassify as “passes” instead of “failures.” Thus any estimate reliant solely on this surrogate would underestimate the failure probability.
@@
In some situations, surrogate predictions may be improved through adaptations to the surrogate model itself. Traditional GPs are limited by the commonly assumed stationarity of the covariance kernel; they are not able to accurately model non-stationary surfaces characterized by regime shifts and stark transitions. Deep Gaussian processes (Damianou and Lawrence,, 2013) are a promising more flexible alternative. They leverage functional compositions of GP layers in which latent layers act as warped versions of the original inputs, allowing for greater flexibility. DGPs have consistently outperformed stationary GPs in modeling non-stationary computer experiments (e.g., Rajaram et al.,, 2021; Yazdi,, 2022; Ming et al.,, 2023), particularly when deployed in a Bayesian framework with full UQ (e.g., Sauer et al.,, 2023b, a). See Sauer, (2023) for a thorough review of DGP surrogates. Crucially, our method is agnostic to surrogate choice as long as UQ is provided. We will demonstrate functionality with both traditional GP and Bayesian DGP surrogates in Section 3.2.
Contour location.
To facilitate failure probability estimation, a surrogate’s sole task it to accurately classify passes and failures. Strategic sequential designs targeting the failure contour, i.e., “contour location,” will outperform space-filling counterparts in this regard. In CL, after initializing with a small space-filling design, subsequent training locations are selected iteratively through the optimization of an acquisition criterion. The computer simulation is evaluated at the selected input(s), the surrogate is updated, and the process is repeated until the budget is exhausted or a stopping criterion is met.

### v2

# Two-stage Design for Failure Probability Estimation with Gaussian Process Surrogates

@@ Abstract
We tackle the problem of quantifying failure probabilities for expensive deterministic computer experiments with stochastic inputs. The computational cost of the computer simulation prohibits direct Monte Carlo (MC) and necessitates a surrogate model, turning the problem into a two-stage enterprise (surrogate training followed by probability estimation). Limited budgets create a design problem: how should expensive evaluations be allocated between and within the training and estimation stages? One may use the entire evaluation budget to sequentially train the surrogate through contour location (CL), with failure probabilities then estimated solely from the surrogate (we call it “surrogate MC”). But extended CL offers diminishing returns, and surrogate MC relies too stringently on surrogate accuracy. Alternatively, a partially trained surrogate may inform importance sampling, but this can provide erroneous results when budgets are limited. Instead we propose a two-stage design: starting with sequential CL, halting CL once learning has plateaued, then greedily allocating the remaining budget to MC samples with high classification entropy. Ultimately, we employ a “hybrid MC” estimator which leverages the trained surrogate in conjunction with the true responses observed in this second stage. Our unique two-stage design strikes an appropriate balance between exploring and exploiting, and outperforms alternatives, including both of the aforementioned approaches, on a variety of benchmark exercises. With these tools, we are able to effectively estimate small failure probabilities with only hundreds of simulator evaluations, showcasing functionality with both shallow and deep Gaussian process surrogates, and deploying our method on a simulation of fluid flow around an airfoil.
Keywords: computer experiment, contour location, deep Gaussian process, emulator, entropy, importance sampling, Monte Carlo
@@ 1 Introduction
Quantifying the probability of failure in complex dynamic systems is of great interest but is challenging due to the computational expense of the so-called “computer experiment.” Consider a deterministic black-box computer simulation, $f:\mathcal{X}\rightarrow\mathbb{R}$ with inputs $\mathcal{X}\subset\mathbb{R}^{d}$ governed by known measure $\mathbb{P}_{\mathbf{x}}$ (with density $p(\mathbf{x})$). The response is thresholded such that $f(\mathbf{x})>t$ indicates failure. The research objective is to estimate the failure probability,
@@
We are particularly motivated by aeronautic applications where system failures may incur immense financial or environmental costs, or even loss of life. In aircraft design and development, computer simulation experiments are a natural, often mandatory, precursor to direct experimentation and prototyping. For example, $f(\cdot)$ may represent the vibrations experienced by an aircraft which become unsafe above certain levels, or the fuel efficiency of an aircraft, which if too low, can be detrimental to the environment, obliterate fuel budgets, or overly restrict travel distances. Inputs $\mathcal{X}$ may encapsulate design parameters and operating conditions (such as wing shapes, flight speeds, and angles of attack). Yet we acknowledge a slew of other relevant applications where quantifying failure probabilities for complex simulations is critical, including structural reliability analysis using finite element simulations (e.g., Mia et al.,, 2017), Boltzmann modeling of chemical processes in fuel filtration (e.g., Belot et al.,, 2021), and thermodynamic flows for nuclear reactors (e.g., Batet et al.,, 2014).
@@
A rudimentary approach to solve Eq. (1) involves brute force Monte Carlo (MC): simply sample an exhaustively large set of inputs, evaluate the computer simulation at each of these configurations, and report the proportion that resulted in failure. But this approach is prohibitively expensive if the computer simulation is costly (incurring long wall clock times), the probability of failure is rare (requiring enormous sample sizes), or both! In this work, we propose a new strategy for quantifying failure probabilities of complex computer simulations when the computation budget restricts evaluations of the simulator to only a few hundred, rendering traditional MC impractical.
@@
In the face of expensive data, statistical models ($\hat{f}$, called “surrogates” or “emulators”) are essential. Surrogates are trained on a limited set of simulator evaluations in order to provide predictions with appropriate uncertainty quantification (UQ) at unobserved inputs. Gaussian processes (GPs) are the canonical choice (Santner et al.,, 2018; Gramacy,, 2020), but deep Gaussian processes (DGPs; Damianou and Lawrence,, 2013) are becoming increasingly popular, particularly for nonstationary computer experiments (Sauer,, 2023). Notably, our contribution is agnostic to surrogate choice as long as effective UQ is provided, and is thus applicable to traditional GPs and Bayesian DGPs (Sauer et al.,, 2023b) and, potentially, Bayesian neural networks (MacKay,, 1995).
Choosing a surrogate model, like a GP or DGP, is only a first step. The utility of a surrogate boils down to two things: how it is trained and how it is used. For our objective (1), surrogate training should prioritize accurate classification of failures and surrogate deployment should focus on effective estimation of $\alpha$. The prevailing question is then: how can a limited budget of expensive simulator evaluations be optimally allocated between and within the training and estimation stages?
@@
We briefly digress to acknowledge some early methods of failure probability estimation, namely the first and second order reliability methods (FORM/SORM; Haldar and Mahadevan,, 1995), which leveraged simplistic linear approximations of the complex model and placed all failure probability estimation in the hands of the “most probable point” (MPP). Although advancements to FORM and SORM have been made with more flexible GP models (e.g., Zhang et al.,, 2015; Su et al.,, 2017), sole reliance on a MPP is too restricting for our target applications.
@@
To solve Eq. (1), surrogate accuracy in classifying failures is of primary importance. Not all training designs are created equally. Space-filling designs (Joseph,, 2016) may miss failure regions altogether, particularly when training budgets are limited. Instead, it is common to train a surrogate on a small initial design then pair it with a relevant acquisition function to select subsequent inputs for evaluation through the simulator. This iterative process of acquisition and surrogate updating can strategically hone in on the contour, all while minimizing evaluations of the expensive simulator. So-called “contour location” (CL) naturally outperforms space-filling counterparts. Acquisition functions seek to balance exploitation and exploration by leveraging both predicted distance from the contour and posterior uncertainty. Popular variations fall under expected improvement frameworks (Bichon et al.,, 2008; Ranjan et al.,, 2008), stepwise uncertainty reduction frameworks (Bect et al.,, 2012; Chevalier et al.,, 2014; Azzimonti et al.,, 2021), or both (Duhamel et al.,, 2023). Others utilize pass/fail classification entropy (Marques et al.,, 2018; Cole et al.,, 2023). Booth et al., (2024b) recently extended CL sequential designs to DGP surrogates. For small failure regions, contour location may be replaced with subset simulation (Au and Beck,, 2001) which initially targets a conservative threshold ($t^{\star}<t$) and iteratively refines it until the true threshold is found ($t^{\star}\rightarrow t$) (Dubourg et al.,, 2011; Huang et al.,, 2016; Hristov et al.,, 2019). Yet these approaches require Markov chain Monte Carlo (MCMC) sampling of cascading input distributions, which we intentionally avoid here.
A well trained surrogate may be leveraged directly within an MC estimator, with surrogate predictions (which are of neglible cost) in place of true simulator evaluations. One could reasonably use the entire budget of simulator evaluations for contour location, then deploy the trained surrogate in such a “surrogate MC” estimator. Although this is common (e.g., Zhu and Du,, 2016; Cheng and Lu,, 2020; Li et al.,, 2021; Lu et al.,, 2023), we find that surrogate learning of failure contours offers quickly diminishing returns. Furthermore, surrogate MC estimators place too much stock in the accuracy of the surrogate and funnel surrogate inaccuracies directly into the failure probability estimate.
Another common approach splits the budget of simulator evaluations into two disjoint stages: some evaluations are used to train the surrogate and some are used for importance sampling (IS; Srinivasan,, 2002; Tokdar and Kass,, 2010). IS hinges on the construction of a bias distribution that intentionally focuses density in the failure region. Inputs are sampled from the bias distribution and are then utilized in a weighted average calculation that corrects for the disparity between the bias density and the original input density. Identifying an effective bias distribution in the face of limited training data is a hefty task (Tabandeh et al.,, 2022). Adaptive methods abound (e.g., Oh and Berger,, 1992; Kurtz and Song,, 2013; Dalbey and Swiler,, 2014; Bugallo et al.,, 2017; Cheng et al.,, 2023). Peherstorfer et al., (2016) propose what they termed “multifidelity importance sampling,” although we prefer the term “surrogate-informed importance sampling” (SIIS) to avoid confusion with multifidelity surrogate modeling (e.g., Park et al.,, 2017). In SIIS, a trained surrogate is used to identify predicted failures, and a bias distribution is fit to these locations (more on this in Section 2.3). SIIS has been deployed with GP surrogates in a variety of subsequent works (e.g., Cole et al.,, 2023; Renganathan et al.,, 2023; Booth et al.,, 2024a; Renganathan,, 2024), and we view it as our primary competitor. Yet SIIS has several crucial drawbacks. We will delay a thorough discussion to Section 2.3, but the highlights are: (i) SIIS relies too heavily on an arbitrary bias distribution, (ii) it largely ignores the carefully trained surrogate, and (iii) it offers no guidance for allocating the budget between surrogate training and IS estimation.
We prefer to avoid a bias distribution altogether, instead focusing on how we could optimally leverage a surrogate and an MC estimator in tandem. The answer lies in a two-stage design. First, we leverage sequential contour location to pin down the failure contour. At the start, it is essential to target inputs which provide the most insight into the entire failure contour (getting the most “bang for our buck” from early observations). As training progresses, successive acquisitions will offer less additional insight, indicated by apparent convergence of the surrogate MC estimator. Although we can not be sure of theoretical convergence with our limited budget, we propose a decision rule to recognize practical convergence (when surrogate learning from CL has saturated). We then spend the remaining simulation budget precisely on the Monte Carlo samples with the highest classification entropy, ultimately training the surrogate on observations from both stages. Our final failure probability estimate is a “hybrid MC” estimator which combines the observations from this second stage with surrogate predictions at unobserved samples.
This manuscript is laid out as follows. Section 2, reviews GP/DGP surrogates, contour location, and importance sampling. Section 3.2 presents our two-stage design and failure probability estimator. Section 4 presents a variety of synthetic benchmarks. Finally, we deploy our method on a computer experiment of an RAE-2822 airfoil in Section 5 and conclude in Section 6.
2 Preliminaries
@@
This estimator is ineffective unless a sufficient number of failures are observed, necessitating very large $M$ particularly for small $\alpha$. With an expensive computer model, we simply cannot evaluate $f(\mathbf{x})$ thousands/millions/billions of times. Throughout, let $N$ denote the total budget of expensive simulator evaluations alotted. We will describe our sample efficient method for estimating failure probabilities in Section 3.2, targeting $N$ in the hundreds, after a quick review of preliminaries.
@@ 2.1 Gaussian process surrogates
A surrogate model $\hat{f}$, trained on a limited budget of simulations, may provide predictions at unobserved inputs in place of the true $f$. To enable our proposed methodology, the surrogate must also provide posterior predictive distributions. Gaussian processes (GPs) are the predominant choice; they assume a multivariate normal distribution over the response. For any $n$ observed locations, our GP prior assumption is $\mathbf{y}_{n}\sim\mathcal{N}(\mu,\Sigma(X_{n}))$. The prior mean $\mu$ is often simplified to the zero vector after centering (which we will use moving forward), and the prior covariance $\Sigma(X_{n})$ is typically a function of inverse euclidean distances, e.g., $\Sigma(X_{n})^{ij}=\Sigma(\mathbf{x}_{i},\mathbf{x}_{j})=k(||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2})$ with kernel $k(\cdot)$. See Santner et al., (2018); Rasmussen et al., (2006); Gramacy, (2020) for thorough reviews. Conditioned on training data $\left\{X_{n},\mathbf{y}_{n}\right\}$, posterior predictions at singleton input $\mathbf{x}$ follow
$\hat{f}_{n}(\mathbf{x})\sim\mathcal{N}\left(\mu_{n}(\mathbf{x}),\sigma_{n}^{2}(\mathbf{x})\right)\quad\textrm{for}\quad\begin{array}[]{rl}\mu_{n}(\mathbf{x})&=\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\mathbf{y}_{n}\\ \sigma_{n}^{2}(\mathbf{x})&=\Sigma(\mathbf{x},\mathbf{x})-\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\Sigma(\mathbf{x},X_{n})^{\top},\end{array}$ (3)
@@
A natural alternative to true MC (2) is “surrogate MC” which leverages surrogate predictions in place of the true simulator,
[…]
Here we use the posterior mean of our GP in place of $f$. If the surrogate is not accurate near the contour, it will misclassify inputs which will severely hinder the effectiveness of this estimator.
Figure 1: Left/Center: Herbie function with failure contour (red) and input distribution (white). Right: Entropy surface (yellow/high, purple/low) for a GP trained on a 100-point LHS (white circles).
Figure 2: Slice of Herbie function along $x_{2}=1.1$. GP was trained on 100-point LHS in 2 dimensions.
For example, consider the two-dimensional “Herbie” function (Lee et al.,, 2011) shown in the left and center panels of Figure 1 (we will revisit the right panel momentarily). We define a failure threshold of $t=1.065$ which demarcates four disjoint failure regions (red contours). We also define $p(\mathbf{x})$ using independent normal distributions centered at the origin (white contours). [^1] Full details of all simulated functions, thresholds, and distributions are provided in the Supplementary Material. We trained a GP surrogate using a random Latin hypercube sample (LHS; McKay et al.,, 2000) of size $n=100$ in two dimensions. Figure 2 shows the GP posterior prediction along the slice $x_{2}=1.1$. The posterior mean (dashed blue) and 95% confidence interval (blue shading) capture the true nonlinear surface (solid black) fairly well. Even so, there are inputs (highlighted by the yellow shading) which the surrogate would misclassify as “passes” instead of “failures.” Thus any estimate reliant solely on this surrogate would underestimate the failure probability.
@@
In some situations, surrogate predictions may be improved through adaptations to the surrogate model itself. Traditional GPs are limited by the commonly assumed stationarity of the covariance kernel; they are not able to accurately model nonstationary surfaces characterized by regime shifts and stark transitions. Deep Gaussian processes (Damianou and Lawrence,, 2013) are a promising more flexible alternative. They leverage functional compositions of GP layers in which latent layers act as warped versions of the original inputs, allowing for greater flexibility. DGPs have consistently outperformed stationary GPs in modeling nonstationary computer experiments (e.g., Rajaram et al.,, 2021; Yazdi,, 2022; Ming et al.,, 2023), particularly when deployed in a Bayesian framework using MCMC sampling to achieve full posterior integration of latent layers (e.g., Sauer et al.,, 2023b, a). See Sauer, (2023) for a thorough review of DGP surrogates. Crucially, our method is agnostic to surrogate choice as long as UQ is provided. We will demonstrate functionality with both traditional GP and Bayesian DGP surrogates in Section 4.
2.2 Contour location
To facilitate failure probability estimation (4), a surrogate must accurately classify passes and failures. Strategic sequential designs targeting the failure contour, i.e., “contour location,” will outperform space-filling counterparts in this regard. In CL, after initializing with a small space-filling design, subsequent training locations are selected iteratively through the optimization of an acquisition criterion. The computer simulation is evaluated at the selected input(s), the surrogate is updated, and the process is repeated until the budget is exhausted or a stopping criterion is met.